OpenStudy (kionnalewis):

what is 500 to the 3 power

OpenStudy (mouses):

125000000

OpenStudy (xxdragoonxx):

500*500*500 =125000000 500^3

OpenStudy (mathmale):

what is 500 to the 3 power?\[500^3=5^3*100^3=?\]

OpenStudy (mathmale):

Note that 100^3=(10^2)^3=10^6

OpenStudy (mathmale):

Welcome to OpenStudy!